OpenStudy (anonymous):

how do i simplify (8+√5)(8-√5) 64-8√5+8√5-√25

OpenStudy (precal):

this is a difference of squares so your middle term will cancel out

OpenStudy (precal):

\[8^2-\sqrt{5}^2\]

OpenStudy (precal):

64-5

OpenStudy (jdoe0001):

for the 1st one, keep in mind as precal said \(\bf \large (a-b)(a+b) = a^2-b^2\)

OpenStudy (precal):

59 is the solution

OpenStudy (precal):

look for these, these are the easy ones to do.....

OpenStudy (anonymous):

so it would be 59 because \[64-\sqrt{25}\] \[64-5=59\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

what happened to the radical 25 @linb

OpenStudy (jdoe0001):

\(\bf 64\cancel{ -8\sqrt{5} }\cancel{ +8\sqrt{5} }-\sqrt{25} \\ \quad \\ 64-\sqrt{5^2}\implies 64-5\implies 59\)

OpenStudy (jdoe0001):

\(\bf \large \sqrt[{\color{red}{ 2}}]{5^{\color{red}{ 2}}}\iff 5\)

OpenStudy (anonymous):

ohhh

OpenStudy (precal):

good job jdoe001

OpenStudy (anonymous):

thanks guys

OpenStudy (precal):

anytime.....

OpenStudy (anonymous):

np